# Towards Reliable Evaluation of Adversarial Robustness for Spiking Neural Networks

### v1

@@ 5.1 Experimental Setup
We conduct experiments on both static datasets (CIFAR-10 and CIFAR-100 [krizhevsky2009learning]) and a dynamic dataset (CIFAR10-DVS [li2017cifar10]). For each dataset, we employ multiple adversarial training methods to obtain robust models and evaluate the attack success rates (ASR) of the proposed approach. Unlike previous studies [bu2023rate, hao2023threaten, lun2025towards], we do not consider non-robust SNNs, as these models often achieve 100% ASR when using APGD with large iteration steps and basic gradient approximation methods.
We train robust SNN models using AT, RAT, AT combined with SR (AT+SR), and TRADES. For robust training, we generate adversarial samples with PGD using 5 iterations and a perturbation magnitude $\varepsilon=8/255$, which is a common practice in current SNN research. The default neuron model in our SNN is the LIF neuron, as described in Eq. 5-Eq. 6, with a time step of 4 for CIFAR-10/100 and 8 for CIFAR10-DVS. For CIFAR-10/100, the network structure is SEWResNet19 [fang2021deep], while for CIFAR10-DVS, we use VGG9 [duan2022temporal]. The default input encoding is direct encoding. The detailed experimental setup and parameter configuration are provided in the Supplementary Materials.
@@ 5.2 Comparison with State-of-the-art Works
Table 2: ASR on different network architectures and depths.
Datasets Architectures Clean Attack BPTR RGA PDSG HART STBP ASSG
CIFAR -10 VGG9 82.09 APGD 62.86 67.15 67.60 77.39 72.37 83.97
SA-PGD 62.86 67.59 67.04 79.86 72.21 88.16
SEWResNet19 78.69 APGD 66.45 63.94 69.31 77.22 75.38 84.06
SA-PGD 66.85 65.08 68.76 79.49 75.11 88.44
SEWResNet31 77.50 APGD 65.72 60.82 68.38 75.73 74.49 83.18
SA-PGD 65.72 61.63 67.27 77.90 74.31 87.38
SEWResNet43 74.80 APGD 57.86 59.25 66.15 75.61 72.79 88.05
SA-PGD 56.87 59.08 64.73 77.41 71.76 91.43
CIFAR -100 VGG9 54.74 APGD 77.28 82.16 81.10 88.06 84.26 89.94
SA-PGD 77.33 82.16 80.63 88.94 84.28 91.77
SEWResNet19 51.90 APGD 82.83 82.50 84.21 89.88 88.22 91.60
SA-PGD 82.99 83.07 84.00 90.80 88.26 93.19
SEWResNet31 51.02 APGD 82.38 80.49 83.86 88.78 87.60 89.74
SA-PGD 82.68 81.08 83.59 89.59 87.66 91.44
SEWResNet43 49.27 APGD 83.24 79.37 84.72 88.00 87.66 89.00
SA-PGD 83.29 80.22 84.46 88.69 87.82 90.60
For gradient approximation methods, we compare ASSG with basic STBP (which uses surrogate gradients during training for adversarial attacks) [wu2018spatio] and with recent methods BPTR [ding2022snn], RGA [bu2023rate], HART [hao2023threaten], and PDSG [lun2025towards]. Implementations follow the original papers. For attack methods, APGD serves as the main baseline, with a default of 100 attack iterations and perturbation magnitude $\varepsilon=8/255$.
Table 1 shows the ASR for different combinations of surrogate gradient approximation methods and attack optimization algorithms. A higher ASR indicates a more accurate evaluation of adversarial robustness. ASSG consistently achieves the highest ASR across various datasets and training methods, particularly on CIFAR-10 and CIFAR10-DVS, where ASSG outperforms other gradient approximation methods by a significant margin (near 10% in 7 cases). When combined with SA-PGD, ASSG further improves ASR over APGD in every experiment. However, SA-PGD shows little benefit when paired with low-quality gradient approximations (BPTR, RGA, PDSG, STBP), indicating that the optimization algorithm and the gradient approximation method mutually influence the effectiveness of adversarial attacks.
Furthermore, we plot the loss curves of SA-PGD on CIFAR-10 for different gradient approximation methods in Fig. 3. As the number of iterations increases, the loss curves of all gradient approximation methods except ASSG tend to converge, while ASSG continues to exhibit an upward trend, further demonstrating its superiority. The loss curves of STBP and HART are higher than that of ASSG in the early stages but are later surpassed by ASSG as the iterations proceed, demonstrating that ASSG dynamically adjusts the spatial–temporal distribution of sharpness parameters. In our configuration, the ASR of RGA and PDSG methods are even lower than that of basic STBP, which contradicts the conclusions of their original papers [bu2023rate, lun2025towards]. However, our results are more reliable, as previous works [bu2023rate, hao2023threaten, lun2025towards] only compared ASR using PGD with 10 iterations. The higher ASR reported in those studies may simply reflect a faster loss increase during the early attack iterations.
@@
Figure 4: ASR for attacks using Atan surrogate gradients with different values of $\alpha$ and ASSG with different values of $A$.
@@
5.3 Ablation Study
Ablation Study on Gradient Approximation. The surrogate gradient function in ASSG is structurally similar to that of Atan, with the key difference being that ASSG uses the upper bound of gradient vanishing degree, $A$, as a parameter. ASSG allows SNNs to adaptively adjust the surrogate gradient sharpness parameter $\alpha_{i,t}^{l}$ in adversarial attacks, whereas Atan uses a fixed sharpness parameter $\alpha$.
Figure 4 shows the ASR for attacks using Atan surrogate gradients with different values of $\alpha$ and ASSG with different $A$, with the target model trained using AT. When $\alpha$ is set as a fixed hyperparameter, a small $\alpha$ leads to inaccurate gradients, whereas an excessively large $\alpha$ causes gradient vanishing, resulting in a unimodal relationship between ASR and $\alpha$. In contrast, ASSG adaptively computes a spatial-temporally heterogeneous $\alpha_{i,t}^{l}$ for each sample, achieving higher ASR than the manually tuned optimal $\alpha$. This demonstrates that the input-dependent adaptive adjustment of surrogate gradient sharpness across spatial-temporal dimensions in ASSG plays a crucial role in accurately estimating input gradients.
Ablation Study on Adversarial Attack. We use ASSG with $A=0.87$ as the surrogate gradient and compare the ASR of multiple attack methods across different numbers of iterations. The update rule in SA-PGD is similar to Adam [kingma2014adam]—incorporating adaptive step sizes and momentum—but it additionally applies per-step $L_{\infty}$-norm clipping to the update (see Eq. 18). To isolate the effect of this clipping, we construct a baseline attack that uses Adam-style optimization for perturbation updates (see Supplementary Materials for details); we denote this variant Adam-PGD. In addition to Adam-PGD, we include the commonly used attacks MI-FGSM [dong2018boosting], PGD, and APGD as baselines. Each attack is evaluated with multiple iteration numbers to ensure a rigorous comparison of convergence behavior and final ASR.
@@ 5.4 Generalizability across Configurations
As shown in Tab. 2, across all network architectures and depths, both ASSG and SA-PGD effectively improve the ASR, demonstrating the excellent applicability of our method to different network structures.
Furthermore, for SNNs based on the SEWResNet19 architecture, we comprehensively evaluate the generalizability of our proposed method in Tab. 3. Since Poisson encoder is a stochastic input transformation, we use Expectation over Transformation (EOT) to compute input gradients during the attack [athalye2018synthesizing], with the number of attack iterations set to 50. To reduce the estimation bias in EOT, the update of the surrogate gradient sharpness in ASSG is frozen within the EOT loop. In addition, the PSN model differs significantly from conventional neuron models—it has no fixed threshold or reset mechanism—making it incompatible with BPTR and RGA methods.
@@ 6 Conclusion
Moreover, our results reveal that existing studies have overestimated the adversarial robustness of SNNs and underestimated the number of attack iterations required for rigorous robustness evaluation. In current SNN adversarial training, the inner maximization in Eq. 2 is typically approximated using 5–10 steps of PGD with the same surrogate gradients used during training (i.e., STBP). Our findings show that this approximation is far from optimal, severely compromising the effectiveness of adversarial training. Hence, there is a pressing need to rethink how to design efficient and reliable methods for solving the inner maximization problem in SNN adversarial training.
[…]
@@
Let $g(x)=\frac{\alpha}{2\left[1+(\frac{\pi}{2}\alpha x)^{2}\right]}$ and the corresponding gradient-vanishing degree function $G(x)=\int_{-x}^{x}g(t)\,dt,x\geq 0$. Assume that $x$ follows a probability distribution $p(x)$ with finite expectation $\mathbb{E}[x]<+\infty$. Then, for any constant $A\in[0,1]$, the expected gradient-vanishing degree satisfies
$\mathbb{E}_{x\!\sim\!p(x)}[G(x)]\leq A\;\text{if}\;\alpha\leq\frac{2}{\pi\mathbb{E}[x]}\tan\!\!\left(\frac{\pi A}{2}\right)$ (S4)
@@
Using the evenness of $g$, we can rewrite $G$ explicitly as
$G(x)$ $=2\int_{0}^{x}g(t)\,dt$
$=\alpha\int_{0}^{x}\frac{1}{1+\bigl(\tfrac{\pi}{2}\alpha t\bigr)^{2}}\,dt$
$=\frac{2}{\pi}\arctan\!\bigl(\tfrac{\pi}{2}\alpha x\bigr).$
Since $G$ is concave on $0,+\infty)$ by Theorem 1, Jensen’s inequality yields
$\mathbb{E}_{x\sim p(x)}[G(x)]\leq G\bigl(\mathbb{E}[x]\bigr)=\frac{2}{\pi}\arctan\!\Bigl(\tfrac{\pi}{2}\alpha\,\mathbb{E}[x]\Bigr).$ (S5)
We now enforce $G(\mathbb{E}[x])\leq A$:
$\frac{2}{\pi}\arctan\!\Bigl(\tfrac{\pi}{2}\alpha\,\mathbb{E}[x]\Bigr)\;\leq\;A.$
Since $A\in[0,1]$, we have $\tfrac{\pi}{2}A\in[0,\tfrac{\pi}{2}]$, where $\tan$ is strictly increasing and $\arctan$ is its inverse. Thus the above inequality is equivalent to
$\arctan\!\Bigl(\tfrac{\pi}{2}\alpha\,\mathbb{E}[x]\Bigr)\;\leq\;\frac{\pi A}{2}\Longleftrightarrow\tfrac{\pi}{2}\alpha\,\mathbb{E}[x]\;\leq\;\tan\!\Bigl(\frac{\pi A}{2}\Bigr),$
that is,
$\alpha\;\leq\;\frac{2}{\pi\,\mathbb{E}[x]}\tan\!\Bigl(\frac{\pi A}{2}\Bigr).$
Under this condition, we have $G(\mathbb{E}[x])\leq A$, and hence by Eq. S5,
$\mathbb{E}_{x\sim p(x)}[G(x)]\;\leq\;G\bigl(\mathbb{E}[x]\bigr)\;\leq\;A,$
which proves Eq. S4. ∎
@@
LIF-2. We use LIF-2 to denote a variant of the LIF neuron model, which has been adopted in prior works such as [bu2023rate, hao2023threaten, lun2025towards, ding2022snn, liu2024enhancing, ding2024enhancing]. The dynamics of LIF-2 are given as follows:
@@
PSN. The PSN is a neuron model that enables parallel processing along the temporal dimension, and its dynamics differ fundamentally from those of conventional neurons [fang2023parallel]. PSN removes the reset mechanism and treats the firing threshold at each time step as a learnable parameter:
@@
We adopt Temporal Efficient Training (TET) [duan2022temporal] as the loss function and optimize the models using SGD with a momentum of 0.9. All models are trained for 100 epochs. For SNNs using neuron models other than PSN, we employ a triangular surrogate gradient during training:
@@
For RGA and PDSG, we follow the parameter settings provided in their original papers [bu2023rate, lun2025towards]. For ASSG, we set $\beta_{1}=0.9$ and $\beta_{2}=0.9$ when Poisson encoding is not used. When Poisson encoding is applied, we instead use $\beta_{1}=0.99$ and $\beta_{2}=0.99$. The initial values of $M_{i,t}^{l}$ and $D_{i,t}^{l}$ are set to $M_{0}=1.0$ and $D_{0}=0.0$, respectively, and the relaxation coefficient $\gamma=1.5$. We search for the optimal value of $A$ within the range $[0.82,0.9]$ with a step size of 0.01. Although the ASR values of ASSG reported in Tab. 1-Tab. 3 correspond to the best-performing $A$, Fig. 4 and Fig. S1 demonstrate that the attack performance of ASSG is robust to the choice of $A$.
For SA-PGD, we set $\beta_{m}=0.5$ and $\beta_{v}=0.8$. When attacking SNNs with Poisson encoding, we estimate the gradient using the Expectation-over-Transformation (EOT) technique [athalye2018synthesizing], with 10 EOT samples per iteration.
@@
S4.2 Additional Comparative Experiments between Atan and ASSG
In this section, we present additional results on models trained with various adversarial training methods, reporting the ASR for attacks using Atan surrogate gradients with different values of $\alpha$ and ASSG with different values of $A$. As shown in the Fig. S1, the results across more models further demonstrate that ASSG’s input-dependent, spatio-temporal adaptive adjustment of the sharpness parameter $\alpha_{i,t}^{l}$ is superior to directly searching for a globally optimal fixed sharpness parameter $\alpha$.
S4.3 Computational Cost Comparison
@@
S4.4 Effectiveness of the Relaxation Term

### v2

@@ 5.1 Experimental Setup
We conduct experiments on both static datasets (CIFAR-10 and CIFAR-100 [17]) and a dynamic dataset (CIFAR10-DVS [19]). For each dataset, we employ multiple adversarial training methods to obtain robust models and evaluate the attack success rates (ASR) of the proposed approach. The surrogate function in ASSG is set to the Atan function in Eq. 7 by default, and the optimal $\omega$ is selected using ternary search.
We train robust SNN models using AT, RAT, AT combined with SR (AT+SR), and TRADES. For robust training, we generate adversarial samples with PGD using 5 iterations and a perturbation magnitude $\varepsilon=8/255$, which is a common practice in current SNN research. The default neuron model in our SNN is the LIF neuron, as described in Eq. 5-Eq. 6, with a time step of 4 for CIFAR-10/100 and 8 for CIFAR10-DVS. For CIFAR-10/100, the network structure is SEWResNet19 [13], while for CIFAR10-DVS, we use VGG9 [12]. The default input encoding is direct encoding. The detailed experimental setup and parameter configuration are provided in the Appendix Sec. S3.
@@ 5.2 Comparison with State-of-the-art Works
For gradient approximation methods, we compare ASSG with basic STBP (which uses surrogate gradients during training for adversarial attacks) [32] and with recent methods BPTR [8], RGA [3], HART [15], and PDSG [22]. Implementations follow the original papers. For attack methods, APGD serves as the main baseline, with a default of 100 attack iterations and $\varepsilon=8/255$.
Table 1 shows the ASR for different combinations of surrogate gradient approximation methods and attack optimization algorithms. A higher ASR indicates a more accurate evaluation of adversarial robustness. ASSG consistently achieves the highest ASR across various datasets and training methods, particularly on CIFAR-10 and CIFAR10-DVS, where ASSG outperforms other gradient approximation methods by a significant margin (near 10% in 7 cases). When combined with SA-PGD, ASSG further improves ASR over APGD. However, SA-PGD shows little benefit when paired with low-quality gradient approximations (BPTR, RGA, PDSG, STBP), indicating that the optimization algorithm and the gradient approximation method mutually influence the effectiveness of adversarial attacks.
Furthermore, we plot the loss curves of SA-PGD on CIFAR-10 for different gradient approximation methods in Fig. 4. The loss curves of STBP and HART are higher than that of ASSG in the early stages but are later surpassed by ASSG as the iterations proceed. As the number of iterations increases, the loss curves of all gradient approximation methods except ASSG tend to converge, while ASSG continues to exhibit an upward trend, demonstrating the effectiveness of ASSG in adaptively evolving the distribution of $\alpha_{i,t}^{l}$ throughout the attack.
In our configuration, the ASR of RGA and PDSG methods are even lower than that of basic STBP, which contradicts the conclusions of their original papers [3, 22]. However, our results are more reliable, as previous works [3, 15, 22] only compared ASR using PGD with 10 iterations. The higher ASR reported in those studies may simply reflect a faster loss increase during the early attack iterations.
@@
5.3 Ablation Study
Ablation Study on Gradient Approximation. ASSG uses $\omega$ as a parameter to adaptively compute the sharpness parameter $\alpha_{i,t}^{l}$ with a certain gradient-vanishing degree, based on the distribution of $\left|u_{i,t}^{l}\right|$ across spatial-temporal dimensions during adversarial attacks. To verify the effectiveness of this input-adaptive and spatial-temporally heterogeneous $\alpha_{i,t}^{l}$, we select five representative surrogate function forms—Atan, Gaussian (Gauss), Sigmoid (Sig), Rectangular (Rect), and Triangular (Tri)—and apply ASSG to perform adversarial attacks on AT-trained SNNs.
For comparison, we determine a non-adaptive fixed $\alpha$ for each surrogate gradient using ternary search. Tab. 2 reports the attack results of SA-PGD. ASSG with different surrogate gradients achieves higher ASR than both the fixed-$\alpha$ baselines and the SOTA methods in Tab. 1. This demonstrates that the input-dependent, spatial-temporally adaptive adjustment of surrogate gradient sharpness in ASSG plays a crucial role in accurately estimating input gradients, and that ASSG generalizes well across surrogate gradients sharing a unified form. We also shows the ASR for attacks using Atan surrogate gradients with different values of fixed $\alpha$ and ASSG with different $A$ in Appendix Sec. S4.2.
Ablation Study on Adversarial Attack. We use ASSG with Atan as the surrogate gradient and compare the ASR of multiple attack methods across different numbers of iterations. The update rule in SA-PGD is similar to Adam [16]—incorporating adaptive step sizes and momentum—but it additionally applies per-step $L_{\infty}$-norm clipping to the update (see Eq. 18). To isolate the effect of this clipping, we construct a baseline attack that uses Adam-style optimization for perturbation updates (see Appendix Sec. S3.2 for details); we denote this variant Adam-PGD. Moreover, we include the commonly used attacks MI-FGSM [11], PGD, and APGD as baselines. Each attack is evaluated with multiple iteration numbers to ensure a rigorous comparison of convergence behavior and final ASR.
@@ 5.4 Generalizability across Configurations
In the Appendix Sec. S4.3, we provide attack results on SNNs with different network architectures and depths. Across all network architectures and depths, both ASSG and SA-PGD effectively improve the ASR, demonstrating the excellent applicability of our method to different network structures.
Furthermore, for SNNs based on the SEWResNet19 architecture, we comprehensively evaluate the generalizability of our proposed method in Tab. 3. Since Poisson encoder is a stochastic input transformation, we use Expectation over Transformation (EOT) to compute input gradients during the attack [2], with the number of attack iterations set to 50. To reduce the estimation bias in EOT, the update of the surrogate gradient sharpness in ASSG is frozen within the EOT loop. In addition, the PSN model differs significantly from conventional neuron models—it has no fixed threshold or reset mechanism—making it incompatible with BPTR and RGA methods.
@@ 6 Conclusion
Moreover, our results reveal that existing studies have overestimated the adversarial robustness of SNNs and underestimated the number of attack iterations required for rigorous robustness evaluation. The results for PSN in Tab. 3 indicate that, without ASSG, SNNs may even exhibit false robustness.
In current SNN adversarial training, the inner maximization in Eq. 2 is typically approximated using 5–10 steps of PGD with the same surrogate gradients used during training (i.e., STBP). Our findings show that this approximation is far from optimal, severely compromising the effectiveness of adversarial training. Hence, there is a pressing need to rethink how to design efficient and reliable methods for solving the inner maximization problem in SNN adversarial training.
@@
Let $g(x)=\alpha k(\alpha x)$ a surrogate function, and the corresponding gradient-vanishing degree function $\int_{-x}^{x}\alpha k(\alpha t)\,dt=\int_{-\alpha x}^{\alpha x}k(\alpha t)\,d\alpha t=G(\alpha x),x\geq 0$. Assume that $x$ follows a probability distribution $p(x)$ with finite expectation $\mathbb{E}[x]<+\infty$. Then, for any constant $A\in[0,1]$, the expected gradient-vanishing degree satisfies
$\mathbb{E}_{x\!\sim\!p(x)}[G(\alpha x)]\leq A\;\text{if}\;\alpha\leq\frac{G^{-1}(A)}{\mathbb{E}[x]}$ (S4)
@@
Recall that for $x\geq 0$,
$G(\alpha x)=\int_{-x}^{x}\alpha k(\alpha t)\,dt=\int_{-\alpha x}^{\alpha x}k(u)\,du,$
hence $G:[0,+\infty)\to[0,1]$ is nondecreasing and satisfies $G(0)=0$ (and typically $\lim_{z\to\infty}G(z)=1$). According to Theorem 1, $G$ is concave on $0,+\infty)$.
Let $X\sim p(x)$ be a nonnegative random variable with $\mathbb{E}[X]<+\infty$. Since $G$ is concave, Jensen’s inequality gives
$\mathbb{E}\bigl[G(\alpha X)\bigr]\leq G\bigl(\mathbb{E}[\alpha X]\bigr)=G\bigl(\alpha\,\mathbb{E}[X]\bigr).$
Therefore, for any $A\in[0,1]$, if
$G\bigl(\alpha\,\mathbb{E}[X]\bigr)\leq A,$
then automatically $\mathbb{E}[G(\alpha X)]\leq A$.
Because $G$ is nondecreasing on $0,+\infty)$, the condition $G(\alpha\mathbb{E}[X])\leq A$ is equivalent to
$\alpha\,\mathbb{E}[X]\leq G^{-1}(A),$
where $G^{-1}(A)$ denotes the (generalized) inverse
$G^{-1}(A):=\inf\{z\geq 0:\;G(z)\geq A\}.$
Thus, it suffices that
$\alpha\leq\frac{G^{-1}(A)}{\mathbb{E}[X]}.$
Combining with the Jensen bound yields
$\mathbb{E}_{x\sim p(x)}[G(\alpha x)]\leq A\qquad\text{if}\qquad\alpha\leq\frac{G^{-1}(A)}{\mathbb{E}[x]},$
which proves the corollary. ∎
According to Corollary 1, we present below a series of specific surrogate functions along with their corresponding methods for computing the adaptive sharpness.
(1) Arctangent (Atan) surrogate
$g_{\mathrm{Atan}}(x)=\frac{\alpha}{2\left[1+(\frac{\pi}{2}\alpha x)^{2}\right]}$ (S5)
$\alpha_{i,t}^{l}=\frac{2}{\pi\,\mathbb{E}[\left\lvert u_{i,t}^{l}\right\rvert]}\tan\!\left(\frac{\pi}{2}A\right)$ (S6)
(2) Gaussian (Gauss) surrogate
$g_{\mathrm{Gauss}}(x)=\frac{\alpha}{\sqrt{2\pi}}\exp\!\left(-\frac{(\alpha x)^{2}}{2}\right)$ (S7)
$\alpha_{i,t}^{l}=\Phi^{-1}(\frac{1+A}{2})\frac{1}{\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]}$ (S8)
where $\Phi$ denotes standard Gaussian PDF.
(3) Sigmoid (Sig) surrogate
$g_{\mathrm{Sig}}(x)=\alpha\,\sigma(\alpha x)\bigl(1-\sigma(\alpha x)\bigr)$ (S9)
where $\sigma(u)=\frac{1}{1+e^{-u}}$.
$\alpha_{i,t}^{l}=\ln(\frac{2}{1-A}-1)\frac{1}{\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]}$ (S10)
(4) Rectangular (Rect) surrogate
$g_{\mathrm{Rect}}(x)=\frac{\alpha}{2}\,\mathbf{1}\!\left\{|x|\leq\frac{1}{\alpha}\right\}$ (S11)
$\alpha_{i,t}^{l}=A\frac{1}{\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]}$ (S12)
(5) Triangular (Tri) surrogate
$g_{\mathrm{Tri}}(x)=\alpha\left(1-\alpha|x|\right)\,\mathbf{1}\!\left\{|x|\leq\frac{1}{\alpha}\right\}$ (S13)
$\alpha_{i,t}^{l}=(1-\sqrt{1-A})\frac{1}{\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]}$ (S14)
It can be observed that the adaptive sharpness parameters of different surrogate gradients share a unified form: $\alpha_{i,t}^{l}=\frac{\omega}{\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]}$. For the Atan, Gauss, and Sig functions, the valid range of $\omega$ is $0,+\infty)$. In contrast, for the Rect and Tri functions, the valid range of $\omega$ is $[0,1]$. This is because for $g_{\mathrm{Rect}}$ and $g_{\mathrm{Tri}}$, the degree of gradient vanishing reaches its upper bound of 1 when $|x|>\frac{1}{\alpha}$, which leads to the constraint $\mathbb{E}\left[\left|u_{i,t}^{l}\right|\right]\leq\frac{1}{\alpha}$ when deriving the inverse function. However, this prerequisite cannot be guaranteed for all inputs and neurons during adversarial attacks. Therefore, we unify the range of $\omega$ of all surrogate functions to $0,+\infty)$.
@@
LIF-2. We use LIF-2 to denote a variant of the LIF neuron model, which has been adopted in prior works such as [3, 15, 22, 8, 21, 10]. The dynamics of LIF-2 are given as follows:
@@
PSN. The PSN is a neuron model that enables parallel processing along the temporal dimension, and its dynamics differ fundamentally from those of conventional neurons [14]. PSN removes the reset mechanism and treats the firing threshold at each time step as a learnable parameter:
@@
We adopt Temporal Efficient Training (TET) [12] as the loss function and optimize the models using SGD with a momentum of 0.9. All models are trained for 100 epochs. For SNNs using neuron models other than PSN, we employ a triangular surrogate gradient during training:
@@
For RGA and PDSG, we follow the parameter settings provided in their original papers [3, 22]. For ASSG, we set $\beta_{1}=0.9$ and $\beta_{2}=0.9$ when Poisson encoding is not used. When Poisson encoding is applied, we instead use $\beta_{1}=0.99$ and $\beta_{2}=0.99$. The initial values of $M_{i,t}^{l}$ and $D_{i,t}^{l}$ are set to $M_{0}=1.0$ and $D_{0}=0.0$, respectively, and the relaxation coefficient $\gamma=1.5$. We search for the optimal value of $A$ within the range $[0.82,0.9]$ with a step size of 0.01. Although the ASR values of ASSG reported in Tab. 1-Tab. 3 correspond to the best-performing $A$, LABEL:assg-abla and Fig. S1 demonstrate that the attack performance of ASSG is robust to the choice of $A$.
For SA-PGD, we set $\beta_{m}=0.5$ and $\beta_{v}=0.8$. When attacking SNNs with Poisson encoding, we estimate the gradient using the Expectation-over-Transformation (EOT) technique [2], with 10 EOT samples per iteration.
@@
Table S2: ASR on different network architectures and depths.
Datasets Architectures Clean Attack BPTR RGA PDSG HART STBP ASSG
CIFAR -10 VGG9 82.09 APGD 62.86 67.15 67.60 77.39 72.37 83.97
SA-PGD 62.86 67.59 67.04 79.86 72.21 88.16
SEWResNet19 78.69 APGD 66.45 63.94 69.31 77.22 75.38 84.06
SA-PGD 66.85 65.08 68.76 79.49 75.11 88.44
SEWResNet31 77.50 APGD 65.72 60.82 68.38 75.73 74.49 83.18
SA-PGD 65.72 61.63 67.27 77.90 74.31 87.38
SEWResNet43 74.80 APGD 57.86 59.25 66.15 75.61 72.79 88.05
SA-PGD 56.87 59.08 64.73 77.41 71.76 91.43
CIFAR -100 VGG9 54.74 APGD 77.28 82.16 81.10 88.06 84.26 89.94
SA-PGD 77.33 82.16 80.63 88.94 84.28 91.77
SEWResNet19 51.90 APGD 82.83 82.50 84.21 89.88 88.22 91.60
SA-PGD 82.99 83.07 84.00 90.80 88.26 93.19
SEWResNet31 51.02 APGD 82.38 80.49 83.86 88.78 87.60 89.74
SA-PGD 82.68 81.08 83.59 89.59 87.66 91.44
SEWResNet43 49.27 APGD 83.24 79.37 84.72 88.00 87.66 89.00
SA-PGD 83.29 80.22 84.46 88.69 87.82 90.60
@@
S4.2 Additional Comparative Experiments between Fixed-$\alpha$ and ASSG
In this section, we present additional results on models trained with various adversarial training methods, reporting the ASR for attacks using Atan surrogate gradients with different values of fixed-$\alpha$ and ASSG with different values of $A$ in Eq. S5-Eq. S6. As shown in the Fig. S1, the results across more models further demonstrate that ASSG’s input-dependent, spatio-temporal adaptive adjustment of the sharpness parameter $\alpha_{i,t}^{l}$ is superior to directly searching for a globally optimal fixed sharpness parameter $\alpha$.
S4.3 Results on Different Network Architectures and Depths
Tab. S2 reports the attack success rates of various attack methods on SNNs with different network architectures and depths. All target models are trained with AT, and the training settings are consistent with Sec. S3.1. ASSG demonstrates significantly stronger attack performance across different network architectures and depths, indicating its strong generalization capability with respect to network structures.
S4.4 Computational Cost Comparison
@@
S4.5 Effectiveness of the Relaxation Term